OpenStudy (dmezzullo):

Solve for x: 2 over 5x + 2 = 10 x = 20 x = 24 x = 4 x = 8

OpenStudy (anonymous):

i hate math i wish i could help sorry

OpenStudy (dmezzullo):

its otttay

hartnn (hartnn):

@dmezzullo u always get so much attention, don't u ? so many viewers....

OpenStudy (anonymous):

\[\frac{ 2 }{ 5x } + 2 = 10\]Multiply both sides by 5x.\[\frac{ 2 }{5x }(5x) + 2 = 10(5x)\]\[2 + 2 = 5x + 10\] Can you solve for x?

OpenStudy (anonymous):

its 20!

OpenStudy (anonymous):

haha i know right lol

OpenStudy (dmezzullo):

@hartnn everybody loves me tht's y

OpenStudy (dmezzullo):

tyvm

hartnn (hartnn):

lol, i know !

OpenStudy (anonymous):

Was 20 right?

OpenStudy (jiteshmeghwal9):

\[\LARGE{\frac{2}{5x+2}=\frac{10}{1}}\]by cross multiplication.we have,\[\LARGE{2(1)=10(5x+2)}\]\[\LARGE{2=50x+20}\]now take all the constants at one side & variables & their coefficients at one side. then solve for 'x' :)

OpenStudy (anonymous):

dmm u have like the most veiws lol